OpenStudy (anonymous):

-2/3v = -6 Please help?

OpenStudy (texaschic101):

is the problem : (-2/3)v = -6 or is it -2/(3v) = -6 ??

OpenStudy (anonymous):

(-2/3)v = -6

OpenStudy (texaschic101):

ok...try dividing both sides by -2/3...but remember that when you are dividing 2 negatives, the sign changes to a positive

OpenStudy (anonymous):

I still don't understand.. I can't do negitives/

OpenStudy (anonymous):

A negative divided by a negative = postive

OpenStudy (texaschic101):

-2/3v = -6 --- divide both sides by -2/3 v = (-6) / (-2/3) when dividing fractions, " flip " what you are dividing by, then multiply v = -6 * -3/2 v = 18/2 which reduces to 9 check.. -2/3(9) = -6 - 18/3 = -6 -6 = -6 (correct) OR...you can do it this way.. -2/3v = -6 --- to get rid of the fraction, multiply everything by common denominator of 3 giving us : -2v = (3)(-6) -2v = -18 -- divide by -2 v = 9...because a neg divided by a neg is a positive. any questions ?

OpenStudy (anonymous):

divide both sides by 2/3 to isolate the "v". 2/3 divided by 6 is is 9. So V=9

OpenStudy (anonymous):

Uhm.. OKay..

OpenStudy (anonymous):

That was literally gibberish.

OpenStudy (anonymous):

Can we try a new problem?

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

(3u/5) = -9

OpenStudy (texaschic101):

if you have a fraction and you want to get rid of it, multiply everything in the equation by the common denominaor. In this problem, it is 5. So we will multiply everything by 5. (5)(3u/5) = (5)(-9) -- the 5's on the left side cancel out 3u = -45 -- now we need to divide both sides by 3 u = -45/3 u = - 15 and you can always check your answer (3u/5) = -9 (3(-15) / 5 = -9 -45/5 = -9 -9 = -9 (correct) answer : u = -15

OpenStudy (anonymous):

Okay. I think I get it now.(: